OpenStudy (anonymous):

Someone please help me with this easy rational equation!! I will be forever in your debt

OpenStudy (anonymous):

x/(x+1)<x/(x-1)

OpenStudy (jdoe0001):

hmm how would you solve , say \(\bf \cfrac{x}{x+1}=\cfrac{x}{x-1}\quad ?\)

OpenStudy (anonymous):

I'd have to multiply for the LCM. But I don't know hw to do that, or what's the LCM of (x+1) and x-1

OpenStudy (jdoe0001):

well.... or just "cross-multiply"

OpenStudy (jdoe0001):

well, you'd do the same in the \(\Large <\) case, you cross-multiply

OpenStudy (arthur_ser):

\[\frac{ x }{x+1}<\frac{ x }{ x-1 }\] \[x(x-1)<x(x+1)\] \[x ^{2}-x<x ^{2}+x\] \[-2x<0\] \[x>0\]

OpenStudy (anonymous):

Thank you all so much!!!